The most recent financial statements for Bello Co. are shown here: Income Statement Balance Sheet Sales $ 20,800 Current assets $ 12,080 Debt $ 16,840 Costs 14,320 Fixed assets 34,650 Equity 29,890 Taxable income $ 6,480 Total $ 46,730 Total $ 46,730 Taxes (25%) 1,620 Net income $ 4,860 Assets and costs are proportional to sales. Debt and equity are not. The company maintains a constant 35 percent dividend payout ratio. What is the internal growth rate? (Do not round intermediate calculations and enter your answer as a percent rounded 2 decimal places, e.g., 32.16.)
Answer #1

ROA=net income/total assets

=(4860/46730)=0.10400171

Retention ratio=1-dividend payout ratio

=(1-0.35)=0.65

Internal growth rate=(ROA*Retention ratio)/[1-(ROA*Retention ratio)]

(0.104001711*0.65)/[1-(0.126151832*0.65)]

which is equal to

=7.25%(Approx).

Thanks

Let me know if it's wrong I will verify